OpenStudy (alizandraw):

The pressure of 1 mol of gas is decreased to 0.5 atm at 273 K. What happens to the molar volume of the gas under these conditions? The molar volume is halved. The molar volume is doubled. The molar volume is tripled. The molar volume stays same.

OpenStudy (abmon98):

Pressure and volume, pressure=force/area, if we have a container with the same number of particles but of different particles, the smaller volume containers will have smaller area and more collision increasing pressure so they have an inversely proportional relationship. Pressure=1/V can you make a guess now?

OpenStudy (alizandraw):

so would it be halved?

OpenStudy (abmon98):

if the pressure is 1/2=1/V cross multiplication V=2

OpenStudy (abmon98):

Pressure and volume, pressure=force/area, if we have a container with the same number of particles but of different particles, the smaller volume containers will have smaller area and more collision increasing pressure so they have an inversely proportional relationship. different volume**

OpenStudy (somy):

\(\huge PV=nRT\) \(\huge V= \frac{nRT}{P}\)

OpenStudy (somy):

Looks like condition was STP before which means it was 1atm as pressure which means pressure was halved

OpenStudy (somy):

if P is halved = whats gonna happen to V?